In a shooting game, 76% of the shots hit the target, and 54 shots missed.
Percentage of shots missed = 100% - 76% = 24%
Total Number of shots fired = _______________ ???

Answer by stanbon(75887)

In a shooting game, 76% of the shots hit the target, and 54 shots missed.
Percentage of shots missed = 100% - 76% = 24%
Total Number of shots fired = _______________ ???
==================================
Let "x" be number of shots fired.
x-54 hit the target
---------------------------
Equation:
x-54 = 0.76x
0.24x = 54
x = 54/0.24 = 225
=======================
Cheers,
Stan H.